OpenStudy (mathlegend):

Trig Identity Prove:

OpenStudy (mathlegend):

\[\frac{ cotx+coty }{ 1-cotxcoty }=\frac{ cosxsiny+sinxcosy }{ sinxsiny-cosxcosy }\]

OpenStudy (accessdenied):

What have you tried here so far? :)

OpenStudy (accessdenied):

When I look at the identity, I see some similarities between the two and some differences. One side has only cotangent (a quotient of cosine and sine), and the other has sines and cosines. The structure of both (two binomials divided) suggests that they are closely related already; perhaps just by converting cotangent to sines and cosines the secret can be revealed.

OpenStudy (mathlegend):

This is what I tried...

OpenStudy (mathlegend):

\[sinxsiny(\frac{ cosx }{ sinx }+\frac{ cosy }{ siny }) / sinxsiny(1-\frac{ cosx }{ sinx }*\frac{ cosy }{ siny })\]

OpenStudy (mathlegend):

\[\frac{ cosxsiny+cosysinx }{ sinxsiny-cosxsinycosysinx }\]

OpenStudy (mathlegend):

Isn't this as simplified as it gets though? I mean I see I could factor out sinx and siny... but that would not get me what I want.

OpenStudy (mathlegend):

@AccessDenied now at this part... I am stuck because the only thing I see is factor... yet factoring won't get me what I want.

OpenStudy (solomonzelman):

for your problem (ORIGINAL PROBLEM) the top of the fraction on the right is cos(x-y) and the bottom of that same fraction is sin(x-y)

OpenStudy (solomonzelman):

http://www.ies-math.com/math/java/trig/kahote/kahote.html

OpenStudy (mathlegend):

So this is dealing with differences?

OpenStudy (solomonzelman):

well, that was just my first thought. you would get tan(x-y) for the right side. I don;t think that would get you anywhere though :( Got to go, sorry and bye.

OpenStudy (phi):

I understand this step \[ sinxsiny(\frac{ cosx }{ sinx }+\frac{ cosy }{ siny }) / sinxsiny(1-\frac{ cosx }{ sinx }*\frac{ cosy }{ siny }) \] but I get \[ \frac{ cosx siny+cosy sinx }{ sinxsiny-cosxcosy } \]

OpenStudy (phi):

which matches your right hand side you can simplify that to sum/differences of angles, and then -tan(x+y) but that is not necessary.

OpenStudy (mathlegend):

So what did I do wrong?

OpenStudy (accessdenied):

You may want to retry the step where you multiplied top and bottom by sin x sin y. It seems you did something like missing the sin x sin y in the denominator of cot x cot y = cos x cos y / (sin x sin y) when multiplying. :)

OpenStudy (accessdenied):

Or perhaps treated the cot x cot y altogether as cosine rather than cotangent, unless there was some typo here.

OpenStudy (mathlegend):

When you multiply the denominators you should get... (sinxsiny) on the left side which I completely understand... but on the right how in the world do you end up with two cosines?

OpenStudy (accessdenied):

\( \displaystyle \frac{ \cot x + \cot y }{1 - \cot x \cot y} \) \( \displaystyle \frac{\sin x \sin y \left( \cot x + \cot y \right) } { \sin x \sin y \left( 1 - \cot x \cot y \right) } \)

OpenStudy (accessdenied):

\( \displaystyle \cot x \cot y = \frac{\cos x}{\sin x} \frac{\cos y}{\sin y} \) You agree with this?

OpenStudy (mathlegend):

Yes

OpenStudy (accessdenied):

By distributing through, you would get this: \( \displaystyle \frac{\sin x \sin y \cot x + \sin x \sin y \cot y}{\sin x \sin y - \sin x \sin y \cot x \cot y } \) which if we switch out cotangents for cosine/sine... \( \displaystyle \frac{\sin x \sin y \frac{\cos x}{\sin x} + \sin x \sin y \frac{\cos y}{\sin y}} { \sin x \sin y - \sin x \sin y \frac{\cos x \cos y}{\sin x \sin y}} \)

OpenStudy (accessdenied):

Earlier I said... "convert cotangents to cosine and sines" or something to that effect. Stuff like this usually happens as a result, we can see things that cancel out like the sin x sin y * (cos x cos y)/(sin x sin y) there. :)

OpenStudy (mathlegend):

So why do I get this?

OpenStudy (accessdenied):

You multiply numerator and denominator by sin x sin y to obtain that expression. The end result is cancellations: \( \displaystyle \frac{\cancel{\sin x} \sin y \frac{\cos x}{\cancel{\sin x}} + \sin x \cancel{\sin y} \frac{\cos y}{\cancel{\sin y}}} { \sin x \sin y - \cancel{\sin x \sin y} \frac{\cos x \cos y}{\cancel{ \sin x \sin y }}} \) Or did I understand your question correctly?

OpenStudy (mathlegend):

Right, so they in the numerator you should have... sinycosx + sinxcosy right?

OpenStudy (accessdenied):

Yes.

OpenStudy (mathlegend):

I don't understand the denominator...

OpenStudy (mathlegend):

Shouldn't you get...

OpenStudy (mathlegend):

Originally the denominator is \[sinxsiny (1-\frac{ cosx }{ sinx }~*\frac{ cosy }{ siny })\]

OpenStudy (mathlegend):

RIght?

OpenStudy (accessdenied):

That is correct.

OpenStudy (mathlegend):

So now when we distribute we should get...

OpenStudy (mathlegend):

\[sinxsiny-cosxsiny~*cosysinx\]

OpenStudy (mathlegend):

Right?

OpenStudy (accessdenied):

That would not be correct... If you had... \( \displaystyle A \left( 1 - \frac{B}{A} \right) \) How would you distribute here?

OpenStudy (mathlegend):

A * 1 = A Then for the fraction... wouldn't the A's reduce... A - B

OpenStudy (accessdenied):

Yes. That is the same framework as the first one we did, as A = sin x sin y and B = cos x cos y. \( \displaystyle \color{red}A \left( 1 - \frac{\color{green}B}{\color{red}A} \right) \) \( \displaystyle \color{red}{\sin x \sin y} \left( 1 - \frac{ \color{green}{\cos x \cos y}}{\color{red}{\sin x \sin y}} \right) \) The first one, you say, is A - B. The comparison would be... \( \sin x \sin y - \cos x \cos y \), right?

OpenStudy (mathlegend):

oh... I think I know why I am messing up... you combined both and I have them separated as two separate fractions...

OpenStudy (mathlegend):

I see it now!

OpenStudy (accessdenied):

Ah! Excellent! :D

OpenStudy (mathlegend):

Thank You guys @AccessDenied and @phi ... I died and came back to life.

OpenStudy (accessdenied):

You're welcome! And I should say thanks to @phi as well for coming by while I was in limbo outside the website. lol :P